OpenStudy (anonymous):

Finding values and composition Suppose f(x)=3x-3, g(x)=x+1, h(x)=x^2+2 and k (x)=|x+4| Find the following: f(k(2))

OpenStudy (anonymous):

\[f(x)=3x-3\\ k(x)=|x+4|\] So, \[f(k(x))=3|x+4|-3\] and \[f(k(2))=3|2+4|-3=\cdots\]